# Mean-field stochastic Volterra equations

### v3

@@ 2. Main results: well-posedness and propagation of chaos
Let $T\in(0,\infty)$, $d,m\in\mathbb{N}$, and let $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ be a filtered probability space, which satisfies the usual conditions. Suppose $B=(B_{t})_{t\in[0,T]}$ is an $m$-dimensional Brownian motion with respect to $(\mathcal{F}_{t})_{t\in[0,T]}$. The law of a random variable $X$ is denoted by $\mathcal{L}(X)$ and, for $p\geq 1$, the space of probability measures on $\mathbb{R}^{d}$ with finite $p$-th moments by $\mathcal{P}_{p}(\mathbb{R}^{d})$. Let $C([0,T];\mathbb{R}^{d})$ be the space of continuous functions from $[0,T]$ to $\mathbb{R}^{d}$, equipped with the supremum norm $\|\cdot\|_{\infty}$, and $\mathcal{P}_{p}(C([0,T];\mathbb{R}^{d}))$ be the space of probability measures on $C([0,T];\mathbb{R}^{d})$ with finite $p$-th moment. For $\rho,\tilde{\rho}\in\mathcal{P}_{p}(\mathbb{R}^{d})$, we write $W_{p}(\rho,\tilde{\rho})$ for the $p$-Wasserstein distance between $\rho$ and $\tilde{\rho}$, see [CD18a, Chapter 5] for its definition, and with a slight abuse of notation, for $\rho,\tilde{\rho}\in\mathcal{P}_{p}(C([0,T];\mathbb{R}^{d}))$ we define $p$-Wasserstein distance
@@
Let us briefly recall the concepts of well-posedness, strong solutions and pathwise uniqueness. We use, for measure spaces $\mathcal{X},\mathcal{Y}$ and $p\geq 1$, the notations $L^{p}(\mathcal{X};\mathcal{Y})$ for the space of all $\mathcal{Y}$-valued, measurable, $p$-integrable functions on $\mathcal{X}$ and, for two Banach spaces $\mathcal{X},\mathcal{Y}$, $C(\mathcal{X};\mathcal{Y})$ for the space of all $\mathcal{Y}$-valued, continuous functions on $\mathcal{X}$. An $(\mathcal{F}_{t})_{t\in[0,T]}$-progressively measurable stochastic process $(X_{t})_{t\in[0,T]}$ in $L^{p}(\Omega\times[0,T];\mathbb{R}^{d})$, on the given probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$, is called a (strong) $L^{p}$-solution of the mean-field SVE (2.1) if
@@ 2.1. Mean-field SVEs with Lipschitz continuous coefficients
    $|\mu(t,x,\rho)-\mu(t,\tilde{x},\tilde{\rho})|+|\sigma(t,x,\rho)-\sigma(t,\tilde{x},\tilde{\rho})|\leq C_{\mu,\sigma}\big{(}|x-\tilde{x}|+W_{\delta}(\rho,\tilde{\rho})\big{)},$
@@
Suppose that the initial value $X_{0}$ is in $L^{p}(\Omega;\mathbb{R}^{d})$, the kernels $K_{\mu},K_{\sigma}$ fulfill Assumption 2.1, the coefficients $\mu,\sigma$ fulfill Assumption 2.2, and $p>\max\{\frac{1}{\gamma},1+\frac{2}{\epsilon}\}$, where $\gamma\in(0,\frac{1}{2}]$ and $\epsilon>0$ are given by Assumption 2.1. Then, the mean-field stochastic Volterra equation (2.1) is well-posed in $L^{p}$. Moreover, for any $q\geq 1$, if $X_{0}\in L^{q}(\Omega;\mathbb{R}^{d})$, the unique $L^{p}$-solution $X$ of (2.1) satisfies
@@
The rates of convergence obtained in (2.11) are analogue to the classical rates for ordinary mean-field SDEs with Lipschitz coefficients (see [CD22a, Theorem 3.20]), using $W_{\delta}(\cdots)^{\delta}$ instead of $W_{2}(\cdots)^{2}$ and, consequently, replacing the exponent $2/d$ by $\delta/d$ in (2.11). Note that in the case of ordinary mean-field SDEs one obtains a pathwise propagation of chaos result (meaning that the $\sup$ in (2.7) is inside the expectation operators), which is a stronger type of convergence than the pointwise convergence presented in Theorem 2.4. This weaker type of convergence is caused by the missing availability of a Burkholder–Davis–Gundy inequality for stochastic Volterra processes. However, the rates of convergence provided in Lemma 2.5 seem to be optimal for synchronous coupling methods, since it is shown in [FG15, Theorem 1 and there after] that for terms of the form $\mathbb{E}[W_{\delta}(\bar{\rho}_{N},\rho)^{\delta}]$ the rates in (2.11) are sharp. Consequently, optimality could be only lost in the inequalities (5.17) or (5.18), which, at least in general, appears not to be the case.
@@ 2.2. Mean-field SVEs with Hölder continuous diffusion coefficients
    $|\mu(t,x,\rho)-\mu(t,\tilde{x},\tilde{\rho})|\leq C_{\mu}\big{(}|x-\tilde{x}|+W_{1}(\rho,\tilde{\rho})\big{)},$
@@
First, we establish the well-posedness of the mean-field stochastic Volterra equation (2.12) with Hölder continuous diffusion coefficients. Its proof is based on a Yamada–Watanabe type approach [YW71], which requires essentially a one-dimensional setting and leads to the stronger assumptions on the kernels. Moreover, note that the Hölder continuous diffusion coefficients are required to be independent of the law of the solution, which is essentially a standard assumption for ordinary mean-field stochastic differential equations, cf. [KMBP24].
@@
Suppose Assumption 2.9, and the kernels $K_{\mu},K_{\sigma}$ and the initial condition $X_{0}$ satisfy Assumption 2.7 or Assumption 2.8 with $p$ given as therein. Then, the mean-field stochastic Volterra equation (2.12) is well-posed in $L^{p}$. Moreover, for any $q\geq 1$, if $X_{0}\in L^{q}(\Omega;\mathbb{R}^{d})$, the unique solution $X$ of (2.12) satisfies
@@ 3. On the well-posedness of ordinary stochastic Volterra equations
Let $X$ be the unique ($d$-, $1$- or $N$-dimensional) solution to the stochastic Volterra equation in either of the settings in Lemma 3.1, Lemma 3.2, Lemma 3.3 or Remark 3.4 with $p>\max\{\frac{1}{\gamma},1+\frac{2}{\epsilon}\}$. In the case of Assumption 2.8, we can set $\gamma=\frac{1}{2}$ and $p>2$ given from there. By adapting [PS23b, Lemma 3.1 and Lemma 3.4] to the multi-dimensional setting, it follows that
@@
For a fixed $\rho=(\rho_{t})_{t\in[0,T]}\in C([0,T];\mathcal{P}_{\delta}(\mathbb{R}^{d}))$, the integral equation (4.2) is an ordinary stochastic Volterra equation. Due to Assumption 2.2, the linear growth and Lipschitz condition of Lemma 3.1 are satisfied. Hence, there exists a unique strong $L^{p}$-solution $X^{\rho}=(X^{\rho}_{t})_{t\in[0,T]}$ to the SVE (4.2) and, by Remark 3.5, we get that $\sup_{t\in[0,T]}\mathbb{E}[|X^{\rho}_{t}|^{q}]<\infty$ for all $q\geq 1$ and that the sample paths of $X^{\rho}$ are almost surely continuous. Moreover, note that $\big{(}\mathcal{L}(X^{\rho}_{t})\big{)}_{t\in[0,T]}\in C\big{(}[0,T];\mathcal{P}_{\delta}(\mathbb{R}^{d})\big{)}$, since, by the representation of the Wasserstein distance in terms of random variables (see [CD18a, (5.14)]) and by Remark 3.5, we have
@@
for $t\in[0,T]$, where we used Hölder’s inequality in the drift integral with $\frac{4+2\epsilon}{4+\epsilon}<1+\epsilon$ such that by the choice of $\delta$ in (2.2), $\frac{4+\epsilon}{4+2\epsilon}+\frac{1}{\delta}=1$ and in the diffusion integral with $\frac{2+\epsilon}{2}$ such that (2.3) holds, Burkholder–Davis–Gundy’s inequality, Fubini’s theorem, the integrability of the kernels from Assumption 2.1 and the Lipschitz continuity of $\mu$ and $\sigma$ from Assumption 2.2. Since we have that
@@
(4.6) $1\displaystyle 0\leq\psi_{n}(x)\leq\frac{2}{n|x|^{1+2\xi}},\quad\forall x\in\mathbb{R},\quad\text{and}\quad\int_{a_{n}}^{a_{n-1}}\psi_{n}(x)\,\mathrm{d}x=$ 1 .
@@
(4.7) =:absent:=:= : $I_{1,t}^{n}+I_{2,t}^{n}+I_{3,t}^{n}+I_{4,t}^{n}+I_{5,t}^{n}.$
@@
which also tends to 00 as $N\to\infty$, by (5) and (5.15). ∎

### v4

@@ 2. Main results: well-posedness and propagation of chaos
[…]
@@
Let us briefly recall the concepts of well-posedness, strong solutions and pathwise uniqueness. We use, for measure spaces $\mathcal{X},\mathcal{Y}$ and $p\geq 1$, the notation $L^{p}=L^{p}(\mathcal{X};\mathcal{Y})$ for the space of all $\mathcal{Y}$-valued, measurable, $p$-integrable functions on $\mathcal{X}$ and, for two Banach spaces $\mathcal{X},\mathcal{Y}$, $C(\mathcal{X};\mathcal{Y})$ for the space of all $\mathcal{Y}$-valued, continuous functions on $\mathcal{X}$. An $(\mathcal{F}_{t})_{t\in[0,T]}$-progressively measurable stochastic process $(X_{t})_{t\in[0,T]}$ in $L^{p}(\Omega\times[0,T];\mathbb{R}^{d})$, on the given probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$, is called a (strong) $L^{p}$-solution of the mean-field SVE (2.1) if
@@ 2.1. Mean-field SVEs with Lipschitz continuous coefficients
    $),|\mu(t,x,\rho)-\mu(t,\tilde{x},\tilde{\rho})|+|\sigma(t,x,\rho)-\sigma(t,\tilde{x},\tilde{\rho})|\leq C_{\mu,\sigma}\big(|x-\tilde{x}|+W_{\delta}(\rho,\tilde{\rho}$$\big)$,
@@
Suppose that the initial value $X_{0}$ is in $L^{p}(\Omega;\mathbb{R}^{d})$, the kernels $K_{\mu},K_{\sigma}$ fulfill Assumption 2.1, the coefficients $\mu,\sigma$ fulfill Assumption 2.2, and $p>\max\{\frac{1}{\gamma},1+\frac{2}{\epsilon}\}$, where $\gamma\in(0,\frac{1}{2}]$ and $\epsilon>0$ are given by Assumption 2.1. Then, the mean-field stochastic Volterra equation (2.1) is well-posed in $L^{p}$. Moreover, for any $q\geq p$, if $X_{0}\in L^{q}(\Omega;\mathbb{R}^{d})$, the unique $L^{p}$-solution $X$ of (2.1) satisfies
@@
The rates of convergence obtained in (2.11) are analogue to the classical rates for ordinary mean-field SDEs with Lipschitz coefficients (see [CD22a, Theorem 3.20]), using $W_{\delta}(\cdots)^{\delta}$ instead of $W_{2}(\cdots)^{2}$ and, consequently, replacing the exponent $2/d$ by $\delta/d$ in (2.11). Note that in the case of ordinary mean-field SDEs one obtains a pathwise propagation of chaos result (meaning that the $\sup$ in (2.7) is inside the expectation operators), which is a stronger type of convergence than the pointwise convergence presented in Theorem 2.4. This weaker type of convergence is caused by the missing availability of the standard Burkholder–Davis–Gundy inequality for the solutions of stochastic Volterra equations since they are, in general, not semimartingales. However, the rates of convergence provided in Lemma 2.5 seem to be optimal for synchronous coupling methods, since it is shown in [FG15, Theorem 1 and there after] that for terms of the form $\mathbb{E}[W_{\delta}(\bar{\rho}_{N},\rho)^{\delta}]$ the rates in (2.11) are sharp. Consequently, optimality could be only lost in the inequalities (5.17) or (5.18), which, at least in general, appears not to be the case.
@@ 2.2. Mean-field SVEs with Hölder continuous diffusion coefficients
    $),|\mu(t,x,\rho)-\mu(t,\tilde{x},\tilde{\rho})|\leq C_{\mu}\big(|x-\tilde{x}|+W_{1}(\rho,\tilde{\rho}$$\big)$,
@@
First, we establish the well-posedness of the mean-field stochastic Volterra equation (2.12) with Hölder continuous diffusion coefficients. Its proof is based on a Yamada–Watanabe type approach [YW71], which requires essentially a one-dimensional setting and leads to the stronger assumptions on the kernels. Moreover, note that the Hölder continuous diffusion coefficients are required to be independent of the law of the solution, which is essentially a standard assumption for ordinary mean-field stochastic differential equations as it appears to be a necessary assumption to implement a Yamada–Watanabe type approach, cf. [KMBP24].
@@
Suppose Assumption 2.9, and the kernels $K_{\mu},K_{\sigma}$ and the initial condition $X_{0}$ satisfy Assumption 2.7 or Assumption 2.8 with $p$ given as therein. Then, the mean-field stochastic Volterra equation (2.12) is well-posed in $L^{p}$. Moreover, for any $q\geq p$, if $X_{0}\in L^{q}(\Omega;\mathbb{R}^{d})$, the unique solution $X$ of (2.12) satisfies
@@ 3. On the well-posedness of ordinary stochastic Volterra equations
Let $X$ be the unique ($d$-, 11- or $N$-dimensional) solution to the stochastic Volterra equation in either of the settings in Lemma 3.1, Lemma 3.2, Lemma 3.3 or Remark 3.4 with $p>\max\{\frac{1}{\gamma},1+\frac{2}{\epsilon}\}$. In the case of Assumption 2.8, we can set $\gamma=\frac{1}{2}$ and $p>2$ given from there. Assuming $X_{0}\in L^{q}$ for $q\geq p$, by adapting [PS23b, Lemma 3.1 and Lemma 3.4] to the multi-dimensional setting, it follows that
@@
For a fixed $\rho=(\rho_{t})_{t\in[0,T]}\in C([0,T];\mathcal{P}_{\delta}(\mathbb{R}^{d}))$, the integral equation (4.2) is an ordinary stochastic Volterra equation. Due to Assumption 2.2, the linear growth and Lipschitz condition of Lemma 3.1 are satisfied. Hence, there exists a unique strong $L^{p}$-solution $X^{\rho}=(X^{\rho}_{t})_{t\in[0,T]}$ to the SVE (4.2) and, by Remark 3.5, we get that $\sup_{t\in[0,T]}\mathbb{E}[|X^{\rho}_{t}|^{q}]<\infty$ for $q\geq q$, provided $X_{0}\in L^{q}$, and that the sample paths of $X^{\rho}$ are almost surely continuous. Moreover, note that $\big(\mathcal{L}(X^{\rho}_{t})\big)_{t\in[0,T]}\in C\big([0,T];\mathcal{P}_{\delta}(\mathbb{R}^{d})\big)$, since, by the representation of the Wasserstein distance in terms of random variables (see [CD18a, (5.14)]) and by Remark 3.5, we have
@@
for $t\in[0,T]$, where we used Hölder’s inequality in the drift integral with $\frac{4+2\epsilon}{4+\epsilon}<1+\epsilon$ (noting that $\frac{4+2\epsilon}{4+\epsilon}$ is the conjugate of $\delta/2$) such that, by the choice of $\delta$ in (2.2), $\frac{4+\epsilon}{4+2\epsilon}+\frac{1}{\delta}=1$ and in the diffusion integral with $\frac{2+\epsilon}{2}$ such that (2.3) holds, Burkholder–Davis–Gundy’s inequality applied to the stochastic processes $(\int_{0}^{r}K_{\sigma}(s,t)\big(\sigma(s,X^{\rho}_{s},\rho_{s})-\sigma(s,X^{\tilde{\rho}}_{s},\tilde{\rho}_{s})\big)\,\mathrm{d}B_{s})_{r\in[0,t]}$, Fubini’s theorem, the integrability of the kernels from Assumption 2.1 and the Lipschitz continuity of $\mu$ and $\sigma$ from Assumption 2.2. Since we have that
@@
(4.6) $0\leq\psi_{n}(x)\leq\frac{2}{n|x|^{1+2\xi}},\quad\forall x\in\mathbb{R},\quad\text{and}\quad\int_{a_{n}}^{a_{n-1}}\psi_{n}(x)\,\mathrm{d}x=1.$
@@
(4.7) $=:$ $I_{1,t}^{n}+I_{2,t}^{n}+I_{3,t}^{n}+I_{4,t}^{n}+I_{5,t}^{n}.$
Let us remark that $H_{t}^{\rho}$ and $H_{t}^{\tilde{\rho}}$ are well-defined stochastic Itô integrals due to Assumption 2.7.
@@
which also tends to 0 as $N\to\infty$, by (5) and (5.15). ∎